Two contractors were working together to finish concrete. One can finish 25 square yards in 4 hours. The other can finish 25 square yards in 5 hours. How much concrete can they finish in 8 hours working together?

Answer by jorel1380(2518)

8/4+8/5=40/20+32/20=72/20
72/20x25=90 sq. yds..